OpenStudy (anonymous):

If f(x)=2x-8 and g(x)=sqrtx-5, what is (f。g)(30)?

OpenStudy (anonymous):

\[f(x)=2x-8\\ g(x)=\sqrt x-5\] So, \[\begin{align*}(f\circ g)(x)&=f(g(x))\\ &=2\left(\sqrt x-5\right)-8\\ &=2\sqrt x-18 \end{align*}\]